A study was conducted to find whether there is any relationship between the weight and blood pressure of an individual. The following set of data was arrived at from a clinical study. The first column represents the serial number and the second and third columns represent the weight and blood pressure of each patient.


Weight (x) 65 82 71 73 63 72 71 77 71 77

Blood Pressure (y)(y) 136 145 123 132 175 165 152 158 117 129


Find the coefficient of correlation for this set of data.


1
Expert's answer
2021-01-11T15:37:44-0500

Soln.

"r= {n \sum xy - \sum x \sum y \over \sqrt {\big (n \sum x^2 - {(\sum x)} ^2\big) \big( n \sum y ^2 - {(\sum y)} ^2 \big) } }"

"\sum x =722, \space \space \sum y =1432, \space \space \sum xy =103202""n=10, \space\space \sum x ^2=52412, \space\space \sum y ^2 =208322"

"r={10(103202) - (722 *1432) \over \sqrt {\big (10*52412 - {722}^2 \big) \big (10*208322 - {1432}^2 \big) }}"

"=-0.19595"